OpenStudy (anonymous):

how do i solve this equation? x^2-16x+48=0

OpenStudy (anonymous):

(x-12)(x-4)=0 ===> x=12 or x=4

OpenStudy (sriram):

factorise it as done by raheen

OpenStudy (anonymous):

so is it basically raheen, its like the previous problems we were doing?

OpenStudy (anonymous):

exactly.

OpenStudy (anonymous):

im stuck at x(2x+7)(3x-1)=0

OpenStudy (anonymous):

how do i get it back the other way?

OpenStudy (anonymous):

multiply the brackets first then mutiply x in the product

OpenStudy (kira_yamato):

You can use the quadratic formula \[x= (-b ± \sqrt{b^2-4ac})/2a\] In this case, you can do something like \[x^2-16x+48=0\] \[x= (-(-16) ± \sqrt{(-16)^2-4(1)(48)})/2(1)\] \[x=(16±\sqrt{64})/2\] \[x=8±4\] \[\lfloor x=12 \rfloor\ OR \lfloor x=4 \rfloor\]